OpenStudy (anonymous):

I am really stuck on this one. Does anyone know how to do this? Rationalize the denominator: 4 7 2−

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ 4 }{ \sqrt{7}-2 }\] Here is tthe equation.

OpenStudy (internettroll11):

multiply top and bottom by{ \sqrt{7}+2 }

OpenStudy (internettroll11):

so your denominator will become 3 and on top you will have 4(sqrt7+2)